OpenStudy (anonymous):

i found that the mean for negative binomial is qr/p but i'm not sure about my varaince i foud something complicated since my professor said that the moment generating funtion is (p^r)/(1-qe^t)^r

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

bek.cud u pls repiy.plz

OpenStudy (anonymous):

I put your link in chat, Idk how to solve this

OpenStudy (anonymous):

basollate plzz help

OpenStudy (lgbasallote):

can i just ask what topic is this? (i.e. algebra, statistics, calculus, etc.)

OpenStudy (anonymous):

statistics

OpenStudy (lgbasallote):

hmm our statistics genius isnt online :/ though i think @Callisto can help you :) i hate statistics sorry :p

OpenStudy (anonymous):

how can i find him\her

OpenStudy (lgbasallote):

she'll be coming here..i tagged her...though she may be away from keyboard right now

OpenStudy (anonymous):

oh my god

OpenStudy (anonymous):

eish she said she ddnt learn it

OpenStudy (anonymous):

guys can u pls find someone else who can help

OpenStudy (anonymous):

hi.can u plz help zarkon

OpenStudy (zarkon):

try this...compute \[E[X(X+1)]\]

OpenStudy (anonymous):

how

OpenStudy (zarkon):

Same way you would compute \[E(X)\]

OpenStudy (anonymous):

any clue

OpenStudy (zarkon):

\[\sum_{x=r}^{\infty}x(x+1){x-1\choose x-r}(1-p)^rp^{x-r}\]

OpenStudy (zarkon):

write \[{x-1\choose x-r}\] using factorials

OpenStudy (zarkon):

\[\sum_{x=r}^{\infty}x(x+1){x-1\choose x-r}(1-p)^rp^{x-r}\] \[\sum_{x=r}^{\infty}x(x+1)\frac{(x-1)!}{(x-1-(x-r))!(x-r)!}(1-p)^rp^{x-r}\]

OpenStudy (zarkon):

\[\sum_{x=r}^{\infty}\frac{(x+1)!}{(x-1-x+r)!(x-r)!}(1-p)^rp^{x-r}\] \[\sum_{x=r}^{\infty}\frac{(x+1)!}{(r-1)!(x-r)!}(1-p)^rp^{x-r}\]

OpenStudy (zarkon):

\[r(r+1)\sum_{x=r}^{\infty}\frac{(x+1)!}{(r+1)!(x-r)!}(1-p)^rp^{x-r}\] \[\frac{r(r+1)}{(1-p)^2}\sum_{x=r}^{\infty}\frac{(x+1)!}{(r+1)!(x-r)!}(1-p)^{r+2}p^{x-r}\]

OpenStudy (zarkon):

\[\frac{r(r+1)}{(1-p)^2}\sum_{x=r}^{\infty}\frac{((x+2)-1)!}{((r+2)-1)!((x+2)-1-((r+2)-1))!}(1-p)^{r+2}p^{x+2-(r+2)}\] \[y=x+2\] \[\frac{r(r+1)}{(1-p)^2}\sum_{y=r+2}^{\infty}{y-1\choose (r+2)-1}(1-p)^{r+2}p^{y-(r+2)}\] \[\frac{r(r+1)}{(1-p)^2}\]

OpenStudy (zarkon):

\[Var(X)=E(X^2)-[E(X)]^2=E(X^2)+E(X)-E(X)-[E(X)]^2\] \[=E(X(X+1))-E(X)-[E(X)]^2\] plug in your known values

OpenStudy (anonymous):

but what if i have to use the moment generating function to find the variance

OpenStudy (zarkon):

oh...well that is much easier

OpenStudy (anonymous):

can u plz show me how

OpenStudy (zarkon):

\[\sigma^2=E(x^2)-E(x)^2\] \[=M''(0)-[M'(0)]^2\]

OpenStudy (anonymous):

my MGF is (p/1-qe^t)^r and the mean i found was qr/p

OpenStudy (anonymous):

but i'm having trouble finding M''(t)

OpenStudy (zarkon):

\[M(t)=\frac{p^r}{(1-qe^t)^r}=p^r(1-qe^{t})^{-r}\] \[M'(t)=p^r(-r)(1-qe^{t})^{-r-1}(-qe^t)\] ok

OpenStudy (zarkon):

\[M'(t)=p^r(-r)(1-qe^{t})^{-r-1}(-qe^t)\] \[M''(t)=p^rrq\frac{d}{dt}[(1-qe^{t})^{-r-1}e^t]\] \[M''(t)=p^rrq[(1-qe^{t})^{-r-1}e^t+(-r-1)(1-qe^t)^{-r-2}(-qe^t)e^t]\]

OpenStudy (anonymous):

bekketso do u get this

OpenStudy (anonymous):

let me try to substitute

OpenStudy (anonymous):

do u get the variance